The Art of Problem Solving has begun selling a cookbook called \What would Euler eat?"
If the price of the cookbook is n dollars (n < 72), then it will sell 720 - 10n copies. What
price will maximize the total revenue we receive for the books?

Answer by richard1234(7193)

Wow, I'm actually an AoPS member (but then there's no book called "What would Euler eat")...

Assuming it sells 720 - 10n books at a price of n dollars, we can say that the revenue is n(720 - 10n) = -10n^2 + 720n. This is a parabola, in which the vertex occurs at the vertex n = 720/(2*10) = $36.